OpenStudy (anonymous):

Could you please explain to me how to factor this? 64-2a^(5) Thank you. ^///^

OpenStudy (anonymous):

factor out the minus two first to write \[-2(a^5-32)\] then you have the difference of two fifth powers

OpenStudy (anonymous):

\[x^5-y^5=(x-y) (x^4+x^3 y+x^2 y^2+x y^3+y^4)\]

OpenStudy (anonymous):

replace x by "a" and y by "2" to get your final answer

OpenStudy (anonymous):

I know the answer is: 2(2-a)(16+8a+4a^2+2a^3+a^4... I'm still confused on how the answer is that...

OpenStudy (anonymous):

i wrote it a little differently, but it is the same. if you factor out the two first instead of minus two you get \[2(32-a^5)\]

OpenStudy (anonymous):

but your job is still to factor the difference of two fifth powers. i wrote the formula above with x and y

OpenStudy (anonymous):

Oh! Thanks! I didn't understand how the 2 was infront of the equation. Thank you ^^

OpenStudy (anonymous):

if you want write \[(x-y) (x^4+x^3 y+x^2 y^2+x y^3+y^4)\] with x = 2 and y = a and you will get the answer

OpenStudy (anonymous):

yw